# DualRAG: A Dual-Process Approach to Integrate Reasoning and Retrieval for Multi-Hop Question Answering

### v1

@@ Abstract
Multi-Hop Question Answering (MHQA) tasks permeate real-world applications, posing challenges in orchestrating multi-step reasoning across diverse knowledge domains. While existing approaches have been improved with iterative retrieval, they still struggle to identify and organize dynamic knowledge. To address this, we propose DualRAG, a synergistic dual-process framework that seamlessly integrates reasoning and retrieval. DualRAG operates through two tightly coupled processes: Reasoning-augmented Querying (RaQ) and progressive Knowledge Aggregation (pKA). They work in concert: as RaQ navigates the reasoning path and generates targeted queries, pKA ensures that newly acquired knowledge is systematically integrated to support coherent reasoning. This creates a virtuous cycle of knowledge enrichment and reasoning refinement. Further, through targeted fine-tuning, DualRAG preserves its sophisticated reasoning and retrieval capabilities in smaller-scale models, demonstrating its versatility and core advantages across different scales. Extensive experiments demonstrate that this dual-process approach substantially improves answer accuracy and coherence, approaching, and in some cases surpassing, the performance achieved with oracle knowledge access. These results establish DualRAG as a robust and efficient solution for complex multi-hop reasoning tasks.
@@
Rong Cheng1, Jinyi Liu1, Yan Zheng1, Fei Ni1, Jiazhen Du2, Hangyu Mao2, Fuzheng Zhang2, Bo Wang1, Jianye Hao1 1Tianjin University, 2Independent Researcher
@@ 2.2 Iterative RAG
Although early RAG methods achieved some success in knowledge retrieval, their limitations in complex reasoning tasks have led to the development of Iterative Retrieval-Augmented Generation (iterative RAG) frameworks. These frameworks enable models to retrieve external knowledge multiple times throughout the reasoning process, gradually constructing a complete reasoning path for complex reasoning tasks. IRCoT Trivedi et al. (2023), Iter-RetGen Shao et al. (2023), In Contenx-RALM Ram et al. (2023) leverage previously generated content from LLMs to trigger retrieval at predefined intervals. FLARE Jiang et al. (2023b), DRAGIN Su et al. (2024) utilize model output logits as signals as trigger retrieval signals. PlanRAG Lee et al. (2024), Plan×$\times$×RAG Verma et al. (2024), GenGround Shi et al. (2024) decompose the original question into sub-questions, retrieve information separately, and synthesize a final answer. SlimPLM Tan et al. (2024), MetaRAG Zhou et al. (2024) generate heuristic answers first, then refine them through retrieval. Self-RAG Asai et al. (2024) introduces special tokens during training, allowing models to control the retrieval directly through these tokens.
@@ 3.2 Fine-Tuning for Compact Models
To mitigate this, we construct a specialized dataset to fine-tune smaller models and enhance their capabilities. Using Qwen2.5-72B-Instruct Team (2024), we apply DualRAG to generate 5,000 complete trajectories from the training sets of HotpotQA Yang et al. (2018), 2WikiMultihopQA Ho et al. (2020), and MuSiQue Trivedi et al. (2022). From these trajectories, we derive targeted training data to enhance three key capabilities. (1) Reasoner: Smaller models struggle to determine when to retrieve. To address this, we use the teacher model’s outputs as supervision, enabling the student model to learn explicit retrieval triggers. (2) Entity Identifier: Smaller models often produce redundant or ineffective queries. To refine this, we employ a cross-encoder model for entity linking, aligning key entities with labeled entities and removing redundancies. (3) Knowledge Summarizer: Smaller models struggle to identify implicit connections between retrieved documents and the question. To address this, we assign summarization labels based on whether a document belongs to the ground truth set.
@@
Capability Count
Reasoner 33,342
Entity Identifier 22,109
Knowledge Summarizer 31,617
Sum 87,068
@@ 4.4 Experimental Results
(5) As illustrated in Figure 3, the retrieval mechanism in DualRAG, triggered when the model identifies the need for additional information, effectively reduces iterative steps thereby minimizing interruptions in the LLM’s reasoning process.
@@ 4.7 Case Study
We conduct case studies and find that DualRAG effectively identifies knowledge needs during reasoning, generates tailored queries, and organizes retrieved knowledge to facilitate high-quality answer generation. Detailed examples and analyses are provided in Appendix E.
@@ Appendix D Metrics Details
Appendix E Case Study

### v3

@@ Abstract
Multi-Hop Question Answering (MHQA) tasks permeate real-world applications, posing challenges in orchestrating multi-step reasoning across diverse knowledge domains. While existing approaches have been improved with iterative retrieval, they still struggle to identify and organize dynamic knowledge. To address this, we propose DualRAG, a synergistic dual-process framework that seamlessly integrates reasoning and retrieval. DualRAG operates through two tightly coupled processes: Reasoning-augmented Querying (RaQ) and progressive Knowledge Aggregation (pKA). They work in concert: as RaQ navigates the reasoning path and generates targeted queries, pKA ensures that newly acquired knowledge is systematically integrated to support coherent reasoning. This creates a virtuous cycle of knowledge enrichment and reasoning refinement. Further, through targeted fine-tuning, DualRAG preserves its sophisticated reasoning and retrieval capabilities in smaller-scale models, demonstrating its versatility and core advantages across different scales. Extensive experiments demonstrate that this dual-process approach substantially improves answer accuracy and coherence, approaching, and in some cases surpassing, the performance achieved with oracle knowledge access. These results establish DualRAG as a robust and efficient solution for complex multi-hop reasoning tasks. We release DualRAG at this https URL.
@@
Rong Cheng1, Jinyi Liu1, Yan Zheng1*, Fei Ni1, Jiazhen Du2, Hangyu Mao2, Fuzheng Zhang2, Bo Wang1, Jianye Hao1* 1College of Intelligence and Computing, Tianjin University, 2Independent Researcher
**footnotetext: Corresponding to: Yan Zheng (yanzheng@tju.edu.cn) and Jianye Hao (jianye.hao@tju.edu.cn)
@@ 2.2 Iterative RAG
[…]
@@ 3.2 Fine-Tuning for Compact Models
To mitigate this, we construct a specialized dataset to fine-tune smaller models and enhance their capabilities. Using Qwen2.5-72B-Instruct Team (2024), we apply DualRAG to generate 5,000 complete reasoning trajectories from the training sets of HotpotQA Yang et al. (2018), 2WikiMultihopQA Ho et al. (2020), and MuSiQue Trivedi et al. (2022). We then filter these trajectories, retaining only those that lead to correct answers. From the filtered data, we derive a targeted initial training dataset aimed at improving three key components: the Reasoner, the Entity Identifier, and the Knowledge Summarizer.
Subsequently, we conduct a detailed analysis of the dataset and perform the following post-processing. (1) Entity Identifier: Even the large teacher model sometimes generates redundant or ineffective queries. To refine these, we employ a cross-encoder model to compute similarity between identified entities, perform entity linking, and remove redundant entities. This reduces unnecessary retrieval operations and lowers context overhead. (2) Knowledge Summarizer: Although the large model produces accurate summaries in most cases, it occasionally omits key information in multi-hop reasoning scenarios due to difficulty in identifying implicit connections between retrieved documents and the question. To remedy this, we compare model-generated summaries with ground-truth documents to ensure latent information is preserved, thereby enhancing knowledge integration integrity and reasoning accuracy.
@@
Capability Raw Processed Count
Reasoner 33,342 0 33,342
Entity Identifier 4,951 17,158 22,109
Knowledge Summarizer 22,058 9,559 31,617
Total 60351 26717 87,068
@@ 4.4 Experimental Results
(5) As illustrated in Figure 3, DualRAG significantly reduces the number of iterative retrieval steps, thereby minimizing interruptions to the LLM’s reasoning process. Notably, our framework minimizes retrievals to the lowest feasible level, demonstrating that DualRAG proactively identifies and fulfills knowledge needs rather than relying on repeated trial-and-error. A detailed analysis is provided in Appendix H.2.
@@ 4.7 Case Study
We conduct case studies and find that DualRAG effectively identifies knowledge needs during reasoning, generates tailored queries, and organizes retrieved knowledge to facilitate high-quality answer generation. Detailed examples and analyses are provided in Appendix F.
@@ 5 Conclusion
Future work
In future work, we plan to incorporate reinforcement learning to further enhance DualRAG’s performance. Reinforcement learning can enable the system to better optimize retrieval and reasoning strategies through continuous feedback, improving adaptability and efficiency in complex tasks.
We also aim to apply DualRAG in AI for Science, where structured retrieval and reasoning are crucial. For example, in biological sciences, frameworks like CellAgent Xiao et al. (2024) highlight the potential of LLMs in automating scientific workflows. Similarly, in scientific domains such as software testing Zheng et al. (2019) and tokamak plasma control Degrave et al. (2022), DualRAG can complement deep reinforcement learning by organizing domain knowledge and guiding exploration.
Acknowledgments
This work is supported by the National Key Research and Development Program of China (Grant No. 2024YFE0210900), the National Natural Science Foundation of China (Grant Nos. 62422605, 92370132) and the Xiaomi Young Talents Program of Xiaomi Foundation.
@@ Appendix D Metrics Details
Exact Match (EM)
$\text{EM}(y,\hat{y})=\begin{cases}1,&\text{if }\hat{y}=y\\ 0,&\text{otherwise}\end{cases}$ (11)
Acc
Acc is more lenient than EM; it considers a match successful as long as the reference answer is a substring of the output.
$\text{Acc}(y,\hat{y})=\begin{cases}1,&\text{if }y\text{ is a substring of }\hat{y}\\ 0,&\text{otherwise}\end{cases}$ (12)
F1
The F1 Score measures the token-level similarity between the output and the reference answer, with a value range of $[0,1]$.
$\text{F1}(y,\hat{y})=\frac{2\times P\times R}{P+R}$ (13)
Where Precision ($P$) and Recall ($R$) are defined as follows, and these two metrics are also commonly used for evaluation:
$P=\frac{|\text{CommonTokens}(y,\hat{y})|}{|\text{Tokens}(\hat{y})|}$ (14)
$R=\frac{|\text{CommonTokens}(y,\hat{y})|}{|\text{Tokens}(y)|}$ (15)
For example, for the question: Who is the mother of the director of the film Polish-Russian War (Film)? the reference answer is "Małgorzata Braunek".
  • If the model outputs Małgorzata Braunek, then $EM=1$, $ACC=1$, $F1=1$.
  • If the model outputs The mother of the director of the film ’Polish-Russian War’ is Małgorzata Braunek., then EMEM=0, ACCACC=1, and $F1$ is between 0 and 1, as the reference answer is part of the output.
$\text{Acc}^{\dagger}$
@@
Appendix E Efficiency Analysis
In our framework, the primary contributors to reasoning latency include: (1) the overhead of large model inference, which is largely determined by the number of generated tokens; (2) the number of retrieval triggers, i.e., the number of iterations performed.
To assess the computational efficiency of DualRAG, we conducted a quantitative analysis as follows.
Token Consumption Across Methods.
Table 7 presents the average output token overhead per question for different methods. We observe that DualRAG maintains a similar token budget to existing iterative retrieval-based approaches, with the average consumption consistently below 1000 tokens.
Method HQA WQA MQA Mean
GenGround 978.9 714.0 893.7 862.2
MetaRAG 482.2 687.5 623.9 597.9
DualRAG 601.3 555.2 637.7 598.1
Table 7: Average output token consumption per question across three datasets.
To better understand where computation is spent within DualRAG, Table 8 reports the token usage of each major module: inference (Infer), evidence identification (EI), and knowledge selection (KS). The sum of these components matches the total cost shown previously.
Module HQA WQA MQA Mean
Infer 196.2 203.1 226.6 208.7
EI 192.5 187.1 239.5 206.4
KS 212.6 165.0 171.6 183.0
Total 601.3 555.2 637.7 598.1
Table 8: Token usage per module within DualRAG across datasets.
Retrieval Trigger Frequency.
As shown in Figure 3 of our main paper, we also analyze the number of retrievals triggered per question across methods. The results indicate that DualRAG significantly reduces retrieval iterations, with the vast majority of questions answered using no more than two retrieval rounds.
DualRAG demonstrates a strong balance between reasoning quality and computational efficiency. By minimizing unnecessary retrievals and managing token consumption across modules, it offers a practical solution for real-world deployment scenarios.
Appendix F Case Study
@@
Appendix G Error Analysis
Although we designed a dual-process architecture to optimize multi-hop reasoning, due to the limitations of large model reasoning capabilities and the inherent flaws in retrieval tools, DualRAG may still fail on certain problems. We manually analyzed the error cases in 100 test samples, and after excluding issues caused by knowledge gaps in the retrieval corpus, we found that the errors primarily fall into the following two categories:
LLM reasoning hallucinations (75%)
In some cases, the model may infer incorrect conclusions based on the documents retrieved.
For example, for the question When did John V, Prince of Anhalt-Zerbst’s father die?
After retrieving the following document: John V of Anhalt-Zerbst (Dessau, 4 September 1504 – Zerbst, 4 February 1551), was a German prince of the House of Ascania and ruler of the principality of Anhalt-Dessau. From 1544, he assumed rule of the re-created principality of Anhalt-Zerbst. John was the second (but eldest surviving) son of Ernest I, Prince of Anhalt-Dessau, by his wife Margarete, daughter of Henry I, Duke of Münsterberg-Oels, and granddaughter of George of Poděbrady, King of Bohemia.
DualRAG correctly identifies Ernest I, Prince of Anhalt-Dessau as John V’s father. However, due to the statement From 1544, he assumed rule of Anhalt-Zerbst," the model incorrectly infers that John V’s father died in the same year he assumed the throne, even though the document does not explicitly provide his father’s death date. This requires further retrieval to confirm this information.
Misunderstanding of the question (25%)
In some cases, the model may misunderstand the intent of the question, particularly when the question contains ambiguous terms (e.g., "country") or implicit references.
For example, for the question: What is the award that the performer of the song Sunday Papers earned?
DualRAG incorrectly interprets the question as asking, "What award did the song Sunday Papers earn?" when, in fact, the question is asking about the awards received by the performer of the song.
Appendix H More Discussions
H.1 The Necessity of Multi-Step Retrieval
To evaluate the effectiveness of multi-turn retrieval, we conducted experiments on the HotpotQA dataset using GPT-4o-mini [^1] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/. Specifically, we compared the following three retrieval strategies:
  • NativeRAG-10: A baseline setup consistent with prior experiments, retrieving up to 10 relevant documents in a single step.
  • NativeRAG-200: A variant that retrieves up to 200 documents in one step to simulate broader retrieval coverage.
  • 2HopRAG: A two-hop approach. In the first step, 10 documents are retrieved using the original query. In the second step, each retrieved document is used as a sub-query to retrieve additional documents, resulting in up to 110 documents overall.
Table 10 presents the comparative results of these strategies.
Method $\text{Acc}^{\dagger}$ EM ACC F1
NativeRAG-10 0.553 0.333 0.475 0.523
NativeRAG-200 0.550 0.333 0.475 0.487
2HopRAG 0.681 0.444 0.603 0.575
Table 10: Performance comparison of different retrieval strategies on HotpotQA using GPT-4o-mini.
Although NativeRAG-200 expands the retrieval scope significantly, it does not lead to noticeable performance improvement over NativeRAG-10. In contrast, the two-hop strategy of 2HopRAG yields substantial gains across all metrics.
This demonstrates that:
  • Multi-hop reasoning often requires the gradual construction of an information chain.
  • One-step retrieval, regardless of breadth, struggles to fully capture the sequential dependencies inherent in such tasks.
These results validate the necessity of multi-step retrieval for complex reasoning tasks. A multi-turn approach not only enhances coverage but also aligns better with the step-wise nature of human-like reasoning.
H.2 Iteration Scaling
Existing methods often treat retrieval as a fixed process, either based on static trigger conditions or by first generating heuristic responses and then deciding whether to initiate further retrievals based on the quality of the output. However, such approaches lack a forward-looking understanding of knowledge requirements. This limitation makes it challenging to dynamically adjust the timing and frequency of retrievals, potentially introducing redundant information or causing delays that disrupt the reasoning process.
To quantify this difference, we analyzed the average number of retrieval iterations required by various IterativeRAG methods, as shown in Figure 3 of our paper. We summarize the key findings in the table below:
Method HQA WQA MQA
IRCoT 3.76 4.19 4.26
MetaRAG 2.93 3.14 3.12
DualRAG 1.64 1.99 2.20
Table 11: Average number of retrieval iterations required by different methods.
To further demonstrate the efficiency of our approach, we conducted experiments on the 2WikiMultiHopQA dataset, systematically varying the maximum number of allowed retrieval iterations. The results are summarized in Table 12 and illustrate that our method (RaQ) can achieve strong performance with minimal retrieval overhead. Notably, only two iterations are sufficient to solve most queries effectively. This conclusion is consistent with the statistical trends shown in Figure 3, which highlights the reasoning depth across different methods.
Max Iterations 0 1 2 3 4 5
F1 Score 0.385 0.556 0.736 0.733 0.733 0.722
Table 12: Impact of limiting the maximum number of retrieval iterations on performance (F1 score) for 2WikiMultiHopQA.
We also extended this analysis to other baseline methods. As shown in Table 13, most competing methods require more iterations to reach comparable performance. For example, only GenGround achieves a similar score to ours after eight iterations, whereas DualRAG reaches that level after just two. Other baselines still show a significant performance gap even after more iterations.
Max Iterations 2 3 5 8
Acc F1 Acc F1 Acc F1 Acc F1
IRCoT 66.9 51.9 78.4 60.7 83.6 67.5 85.5 66.9
GenGround 79.8 67.5 81.3 67.2 78.1 70.3 82.4 74.0
MetaRAG 55.3 56.3 56.3 56.9 59.7 61.2 61.6 60.9
DualRAG 83.5 75.6 83.6 75.9 85.0 77.3 84.3 76.4
Table 13: Performance comparison of different methods under varying maximum iteration constraints.
These findings demonstrate that our framework, by proactively identifying knowledge demands, minimizes the number of retrievals—i.e., interruptions to the reasoning process—to its theoretical minimum. In the datasets used, most questions require 1–2 reasoning hops, though some involve up to 4–5 documents. DualRAG closely aligns retrieval steps with this minimal number, reflecting its ability to avoid unnecessary trial-and-error iterations. Rather than relying on simple iterative optimization, DualRAG offers a more efficient and effective retrieval strategy.
H.3 Enhanced Variant of w/o KO Ablation
To further assess the importance of the Knowledge Outline (KO) mechanism in our framework, we conducted an enhanced variant ablation, denoted as w/o KO.
While the original w/o KO setting removes the KO module entirely and directly feeds raw retrieved documents into the Reasoner, the w/o KO variant retains the summarization step but skips the construction of a structured knowledge outline. This experiment simulates a setting where the model must reason based on flat summaries, without the benefit of cross-document structuring provided by KO.
The results of w/o KO are presented in Table 14, showing a consistent decline in performance across all benchmarks:
Method HQA WQA MQA
Acc F1 Acc F1 Acc F1
DualRAG 70.0 65.7 85.0 77.3 66.3 56.3
- w/o KO 69.5 64.0 79.4 72.8 58.1 49.8
- w/o KO 69.7 65.1 74.3 66.2 55.8 49.2
Table 14: Performance of the enhanced w/o KO variant where the Knowledge Outline is removed, but summarization is retained.
These results demonstrate that while summarization alone does offer some level of abstraction, it falls short of the benefits provided by a structured Knowledge Outline. Without KO, the model struggles to effectively organize and leverage cross-document information, which not only impairs reasoning quality but may also lead to inefficiencies—such as processing overly verbose or less relevant content. This highlights the importance of KO as both a cognitive aid for the LLM and a practical design for scalable reasoning.
H.4 Ablation on the Reranker
To assess the impact of the reranker module in our framework, we conducted an ablation study. Specifically, we modified the original retrieval setup—which first retrieves the top-50 candidates and then applies a reranker to select the top-10 (as described in lines 470–480)—to directly retrieve the top-10 documents without reranking. The comparative results are shown in Table 15.
Method HQA WQA MQA
Acc F1 Acc F1 Acc F1
DualRAG 70.0 65.7 85.0 77.3 66.3 56.3
- w/o rerank 67.6 63.3 86.3 76.4 64.8 54.9
Table 15: Ablation results on the reranker component across three datasets.
As shown, removing the reranker module does lead to a slight drop in performance, though the decrease is not as substantial as that observed when ablating other core components of our system (see Table 3). This observation highlights the robustness of our design.
The reranker serves as a more precise scoring model during retrieval. The relatively modest performance degradation observed in this ablation suggests that the dual-process structure of DualRAG is capable of partially compensating for limitations in retrieval quality. In other words, our design enables the reasoning capabilities of large language models (LLMs) to better utilize—and even enhance—the utility of standard retrieval tools, thereby narrowing the gap with stronger retrieval systems.
We would also like to emphasize two additional points:
  • The reranker is not redundant: It still contributes meaningfully to retrieval quality. However, thanks to our structured and adaptive framework, even with a simpler retriever, comparable results can be achieved.
  • Fair comparison across baselines: All models in our experiments, including DualRAG and the baselines, use the same retrieval settings—including reranker integration where applicable—ensuring a fair and consistent basis for comparison.
H.5 More Baselines
Methods such as FLARE Jiang et al. (2023b) and DRAGIN Su et al. (2024) dynamically decide when to perform retrieval and which tokens to query based on log-probability signals. While sharing similar objectives with our method—namely, improving the timing and targeting of retrieval—there are fundamental differences in implementation. FLARE relies on the model’s output uncertainty to trigger retrieval, whereas our method proactively anticipates and activates retrieval based on predicted knowledge demands, thereby enabling more structured and efficient knowledge usage.
Nonetheless, FLARE represents a strong and relevant baseline. We plan to include a more detailed discussion of such approaches in the revised version of our paper.
To further validate the effectiveness of our method, we conducted additional experiments using GPT-4o-mini. The results are shown in Table 16.
Method HotpotQA 2Wikimultihopqa MuSiQue
FLARE 0.728 0.427 0.661 0.573 0.687 0.504 0.772 0.632 0.665 0.355 0.628 0.502
DualRAG 0.767 0.468 0.679 0.617 0.784 0.612 0.814 0.722 0.659 0.383 0.611 0.528
Table 16: Comparison of DualRAG and FLARE on three MHQA datasets using GPT-4o-mini.
H.6 Other LLMs
We chose the Qwen series models for our experiments based on the following considerations:
  • Strong natural language processing capabilities. The Qwen series models have demonstrated excellent performance across a wide range of natural language processing tasks, with outstanding knowledge integration and reasoning abilities.
  • Open-source reproducibility. The Qwen series models are open-source, and the community provides stable API support, which ensures high reproducibility of our experiments and facilitates further research and validation by the community.
DualRAG does not rely on a specific LLM architecture, as the LLM, serving as the core "brain" of DualRAG, offers high scalability and flexibility, making it not limited to any particular model.
To verify the applicability of our approach across different LLM architectures, we also conducted experiments on other open-source models (Llama-3.1-70B-Instruct [^2] https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct and Mistral-Small-24B-Instruct-2501333https://huggingface.co/mistralai/Mistral-Small-24B-Instruct-2501). The results are shown in Table 17.
Method HotpotQA 2Wikimultihopqa MuSiQue
Base LLM LLaMA-3.1-70B-Instruct
Direct 0.546 0.267 0.364 0.415 0.388 0.254 0.349 0.350 0.306 0.091 0.191 0.199
Native 0.728 0.478 0.534 0.610 0.479 0.370 0.389 0.422 0.408 0.260 0.301 0.342
IRCoT 0.816 0.528 0.710 0.654 0.733 0.557 0.790 0.644 0.606 0.338 0.522 0.457
GenGround 0.835 0.514 0.713 0.665 0.843 0.692 0.862 0.781 0.593 0.409 0.563 0.492
DualRAG 0.834 0.522 0.732 0.675 0.890 0.730 0.887 0.815 0.684 0.405 0.634 0.537
Base LLM Mistral-Small-24B-Instruct-2501
Direct 0.367 0.229 0.246 0.323 0.322 0.277 0.291 0.326 0.161 0.074 0.096 0.148
Native 0.646 0.436 0.475 0.572 0.411 0.315 0.348 0.384 0.367 0.235 0.274 0.339
IRCoT 0.618 0.403 0.584 0.527 0.419 0.310 0.645 0.396 0.288 0.172 0.264 0.251
GenGround 0.726 0.421 0.640 0.572 0.710 0.510 0.765 0.618 0.578 0.298 0.538 0.430
DualRAG 0.744 0.441 0.667 0.604 0.761 0.549 0.785 0.650 0.601 0.310 0.567 0.461
Table 17: Evaluation results of DualRAG and baselines on three MHQA datasets using two different open-source LLMs. Bold indicates best performance.
These experimental results indicate that our method is also applicable to other LLM architectures, demonstrating robust performance across different models.